OpenStudy (anonymous):

solution of 40-2L/10=54-3k/15

OpenStudy (anonymous):

L = k - 70

OpenStudy (anonymous):

Is the problem: (40-2L)/10=(54-3k)/15? If so then the first step is multiply by 10 and 15: 15(40-2L) = 10(54-3k) Distribute: 600-30L = 540-30k Subtract 540 from both sides 60-30L = -30k Divide by 30 2-L = -k or 2+k = L

OpenStudy (anonymous):

You could be correct I took the problem as 40-(2L/10) = (54 - 3k)/15 But I think you are correct TBates.